Design of Algorithms - Course Review

K. Subramani LCSEE, West Virginia University, Morgantown, WV {ksmani@csee.wvu.edu} January 13, 2011

1. Have you me the pre-requisites of this class? (a) Yes, I have taken both CS 220 and CS 221.

(b) No, I have taken CS 220, but not CS 221.

(c) I have taken neither CS 220 nor CS 221.

2. Are you familiar with mathematical induction?

(a) Extremely familiar with both the first and second principle.

(b) Familiar with the first but not the second.

(c) I am unfamiliar with induction.

3. Are you familiar with proof techniques, such as contradiction, contraposition and direct proof?

(a) Yes, very familiar.

(b) Somewhat familiar.

(c) Unfamiliar.

4. Are you familiar with modeling and solving firstorder and divide-and-conquer recurrences?

(a) Yes, very familiar.

(b) Somewhat familiar.

(c) Unfamiliar.

5. Are you familiar with the following algorithm design techniques:

(a) Divide-and-Conquer (Very familiar − Somewhat Familiar − Unfamiliar).

(b) Dynamic Programming (Very familiar − Somewhat Familiar − Unfamiliar).

(c) Greedy Methods (Very familiar − Somewhat

Familiar − Unfamiliar).

6. Does the instructor address your concerns in timely fashion?

(a) Yes.

(b) No.

(c) Most of the time.

7. Do you prefer interactive lectures?

(a) Yes, the instructor should ask me questions.

(b) No, the instructor should leave me alone.

8. Are you comfortable with the revised syllabus? (a) Yes.

(b) No, more material needs to be taken out.

(c) No, more material needs to be added.

9. What is your motivation for taking this course? (a) I am interested in algorithm design.

(b) I need it to graduate.

10. What is the most important outcome for you? (a) I need to get good grades.

(b) I need to learn a lot.

11. What aspects of the course would you like to change and in what way?